OpenStudy (anonymous):

A length of wire has resistance R. The same wire twice as long has resistance __________. A. 2R B. R C. R/2 D. None of the options are correc

OpenStudy (sidsiddhartha):

resistence of a wire of length L area of cross section A and resistivity P is \[R=P \frac{ L }{ A }\] so from this relation we can see that resistence is directly proportional to lenght \[R \alpha L \rightarrow \frac{ R_1 }{ R_2 }=\frac{ L_1 }{ L_2 }\] okay so far?

OpenStudy (anonymous):

ok I follow!!

OpenStudy (sidsiddhartha):

now the wire has lenght twice so \[\frac{ R_1 }{ R_2 }=\frac{ L_1 }{ L_2}=\frac{ L_1 }{ 2L_1 }=\frac{ 1 }{ 2 }\] so \[R_2=2*R_1\] so if u double the lenght then resistence will also be doubled

OpenStudy (anonymous):

so if length increase by double the resistance is 2R

OpenStudy (sidsiddhartha):

yup :)

OpenStudy (anonymous):

Thank you truly!!!

OpenStudy (anonymous):

if I have more questions can I ask ?

OpenStudy (sidsiddhartha):

sorry bud it's getting late here time for me to hit the bed now :)

OpenStudy (sidsiddhartha):

you can ask i can wait a little

OpenStudy (anonymous):

The origin of the glow from the filament inside the ac incandescent lamp in your bedroom is A. electrons resupplied by the wall outlet B. the positive charges that flow in the filament C. resistance in the filament D. electrons already in the filament

OpenStudy (anonymous):

i think its D

OpenStudy (sidsiddhartha):

yeah D

OpenStudy (anonymous):

thank u

OpenStudy (sidsiddhartha):

np :)

OpenStudy (anonymous):

The current in a 100-W bulb connected to a 120-V source is __________. A. 1.2 A B. 12,000 A C. 0.5 A D. None of the options are correct.

OpenStudy (anonymous):

sorry please help !!